All of the answers are correct, I just need to see how to do parts C-H, please show step by step! THANK YOU! I'll be sure to thumbs up!:)

A)   Assume the following:

       Number of college students in the market:                       25,000

       Average number of copies per student per year                     125

       Your company’s estimated share of the total market               13%

SOLVE FOR THE FOLLOWING:

Total annual market demand                                                            3,125,000

Estimated annual company demand                                                     406,250

Estimated monthly demand                                                                    33,854

Estimated weekly demand                                                                        7,813

B)  Given the following for your copy business:   

Annual Expenses:

     VC  Paper:                       $  4,062.50                  FC  Salaries:                              $ 30,000                          

      FC  Rent:                        $ 12,000                      VC  Hourly labor:                      $ 62,400

      FC  Insurance:                $   2,500                       FC  Maintenance contract         $   2,000

      FC Advertising:             $   1,800                        FC Utilities:                             $    2,400

      VC Other Supplies:        $   5,000

     Total Fixed Cost:    $ 50,700                       Total Variable Cost: $71,462.50.                 Fixed Cost/unit: 50700/406250= $ .12    Variable Cost/unit:  71462.5/406250=$ .175  or $ .18

C.  Calculate the selling price per copy if you must have a markup of 25% on the selling price.

                                         $ .40

D. What is the selling price per copy if you desire a 150% markup on cost?

                                          $ .75

E.  What is the markup on selling price with a $ .50 selling price per copy?

                                         40%

F.  What is the markup on cost with a $ .50 selling price?

                                          . 667  or 67%

G.   If you charge $.45 per copy, what is the highest your expenses can be and still maintain a

        25% markup on selling price?

                                               $ .34   

H.   If you charge $.45 per  copy, what is the highest your expenses can be and still

        maintain a 150% markup on costs?

                                              $ .18
Answer #1

Total cost per unit = 0.18+0.12 =0.3

(i) If 25% markup on selling price is needed and selling price is S, then

S*(1-0.25) = cost

S= 0.3/0.75 = 4

(ii) Selling price per copy when 150% markup on cost is needed

S=(1+1.5)*0.3 = 0.75

(iii) If selling price is S,

S*(1-0.4) =0.3

S=0.3/0.6 =0.5

(iv) If the markup on cost is x

(1+x)*0.3 =0.5

1+x= 0.5/0.3

x=0.6666

(v) If the expense is E, then

S*(1-0.25)=E

E = 0.45*0.75 =0.34

(vi) (1+x)*cost = S

(1+1.5)* cost =0.45

cost =0.45/2.5 = 0.18